OpenStudy (anonymous):

use the linear aproximation L(x) of the function f(x)= 5th root of x to the vaule 5th root of 31. find an x interval on which aproximation is with .1 of the actual vaule. use L(x) = f(a)+f'(x)(x-a)

OpenStudy (turingtest):

well, we know the 5th root of 32, so use the approximation near the point x=32

OpenStudy (turingtest):

that is, a=32

OpenStudy (turingtest):

that should be L(x) = f(a)+f'(a)(x-a) btw

OpenStudy (anonymous):

this practice exam has x in there. and why do we use 32 for a?

OpenStudy (turingtest):

linear approximations are done around a known value x=a. we can find the 5th root of 32 easily, so let's use that point since it's near 31

OpenStudy (anonymous):

f'(32) = 1/5 *32?

OpenStudy (turingtest):

no, what is f'(x) ?

OpenStudy (anonymous):

the derivitive of f(x)

OpenStudy (turingtest):

yeah I know, I mean what is the derivative of f(x) in this case?

OpenStudy (anonymous):

\[f \prime(x)=\sqrt[5]{32}\]

OpenStudy (turingtest):

no, that is f(x)

OpenStudy (turingtest):

actually that's f(32) first off, what is f(x) ? then what is the derivative f'(x) ? finally what is f(32) and f'(32) ?

OpenStudy (anonymous):

f(x) is 5th root of x f'(x) wouls be 1/5x^(4/5)

OpenStudy (turingtest):

f'(x)=1/5x^(-4/5)

OpenStudy (turingtest):

so what are f(32) and f'(32) ?

OpenStudy (anonymous):

if you move it to the denominator it becomes positive tho right?

OpenStudy (turingtest):

yes, I didn't know it was supposed to be in the denom

OpenStudy (anonymous):

5th root of 32 is 2 f'(32) is 1/80

OpenStudy (turingtest):

yes, so what is L(x) ?

OpenStudy (anonymous):

L(x)=2+(1/80)(x-32)

OpenStudy (turingtest):

good, so what is L(31) ?

OpenStudy (anonymous):

159/80

OpenStudy (anonymous):

or 1.9875

OpenStudy (turingtest):

and that makes sense, being a little less than 2 so we're done

OpenStudy (anonymous):

awesome that made it much more clear to me, thank you

OpenStudy (turingtest):

welcome!